OpenStudy (anonymous):

What is the slope of a line parallel to 10x - 5y = 8? If you help me get the answer, I'll give out a medal <3

OpenStudy (johnweldon1993):

lol aww a medal!!!? haha :P okay .,..well we want to get this into slope intercept form..how do we do that?

OpenStudy (johnweldon1993):

y = mx + b

OpenStudy (johnweldon1993):

we want to isolate y right? 10x - 5y = 8 subtract 10x from both sides -5y = 8 - 10x divide both sides by -5 y = -8/5 + 2x written correctly y = 2x - 8/5 what is your slope? remember y=mx + b.....m is your slope and y is your y-intercept...so what is your slope?

OpenStudy (anonymous):

@johnweldon1993 b is the y-intercept.

OpenStudy (johnweldon1993):

i meant that...sorry about that typo @jazy

OpenStudy (johnweldon1993):

once you find that your m (slope) is ...... remember that the slope of a parallel line...is the same slope of the original line..so if your slope of THIS line is ..... the slope of the line you want is ALSO .... what is your slope?

OpenStudy (anonymous):

so 2 is the x would 8/5 be the slope? I'm so lost:////

OpenStudy (johnweldon1993):

no no...you have it.... y = mx + b m is your slope...x is just an x...and b is your y - intercept we only want the slope which is what you said...it's 2 y = mx + b y = 2x - 8/5 see the similarity? 2 (which is m...which is the slope) do you understand it better now? 2 is the m....m is the slope...so 2 is the slope

OpenStudy (johnweldon1993):

and then remember that the slopes of parallel lines...are the same...so the slope of the line you want is?